I have a worksheet
The measure of the vertex angle of an isosceles triangle is 126. Find the measure of each base angle.
I'm not sure how to do this.
Also,
If the measure of an exterior angle drawn at vertex M of triangle LMN is x, then m angle L + m angle N is what?
I am so confused. I don't understand this geometry at all.

Answer by stanbon(48535)

The measure of the vertex angle of an isosceles triangle is 126. Find the measure of each base angle.
Draw the picture of an isosceles triangle.
Let the vertex angle be 126 degrees
That leasve 180-126 = 54 degrees for the other two angles combined.
Each of them must be 54/2 = 27 degrees
---------------------------------
Also,
If the measure of an exterior angle drawn at vertex M of triangle LMN is x, then m angle L + m angle N is what?
Draw a triangle.
Draw an exterior angle; it is the supplement of the interior angle where you
drew it. The measure of the exterior angle is x; its supplement is 180-x
-------
Then the measure of the sum of the other two angles is 180-[(180-x)] = x
------------
Cheers,
Stan H.